OpenStudy (anonymous):

Let h(x)= (sqrtx)(f(x)). If f(2) =5 and f'(2) = -3, find h'(2)

OpenStudy (anonymous):

\[g(x)=\sqrt{x}f(x)\] right?

OpenStudy (anonymous):

or rather \(h(x)\) no matter use the product rule to find the derivative, then plug in the numbers

OpenStudy (anonymous):

you got the derivative?

OpenStudy (anonymous):

so it would be (x^1/2)(f(x)) + F'(x)(sqrtx) and I just plug in the numbers it gave me?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the derivative of \(\sqrt{x}\) is not \(x^{\frac{1}{2}}\)

OpenStudy (anonymous):

\(x^{\frac{1}{2}}\) is the same as \(\sqrt{x}\) written in exponential notation (which is no help) you need the derivative of \(\sqrt{x}\), do you know it?

OpenStudy (anonymous):

no I don't can you explain how to get it?

OpenStudy (anonymous):

yes, you memorize it \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

you can if you like rewrite \(\sqrt{x}=x^{\frac{1}{2}}\) and then take the derivative using the power rule it is dumb to do it this way, because the square root if a very common function, so it is much better just to know it like knowing \(9\times 8=72\)

OpenStudy (anonymous):

also math teachers love using it on tests and quizzes while your colleagues are wasting their time using the power rule, go right to the answer you have memorized much quicker and easier

OpenStudy (anonymous):

That made this problem a lot easier to understand thanks I got it from here !

OpenStudy (anonymous):

yw